OpenStudy (anonymous):

Pre-Calc : How do you get the direct relationship between y and x and how do you know if the parametric equations determine y as a function of x with the following? x = 2t and y = 3t -1

OpenStudy (anonymous):

y=1.5x since x=2t and y=3t-1 has 1.5 x's.

OpenStudy (anonymous):

OOPS! y=1.5x-1

OpenStudy (anonymous):

a direct relationship between y and x means that y = kx. this also is equivalent to saying that y varies directly as x , y is proportional to x, and that the graph of y goes through the origin and is linear . for your second question, solve for t,

OpenStudy (anonymous):

x = 2t, so t = x/2, now plug that into y, y = 3 ( x/2) -1. is y now a function of x?

OpenStudy (anonymous):

a direct relationship between y and x means that y = kx for some k not zero, a constant k

OpenStudy (anonymous):

Thank you so much! I had trouble understanding this in school.

OpenStudy (anonymous):

its a special linear function, y = kx , where the y intercept is zero (passes through (0,0)